# On Supportedness in Multi-Objective Combinatorial Optimization

### v1

@@ 3 Supported and Weakly Supported Non-dominated Points in MOCO problems
Several definitions and characterizations of supported non-dominated points exist in the literature.
@@
    if $y^{\prime}$ is non-dominated and $y^{\prime}$ lies on the boundary of the convex hull of the feasible region $\operatorname{conv}(\mathcal{Y})$ in the outcome space, i.e., $y^{\prime}\in\mathcal{Y}_{N}\cap\partial(\operatorname{conv}(\mathcal{Y}))$ (Eusébio and Figueira, 2009c, ; Schulze et al.,, 2019). The set of supported non-dominated points obtained with respect to this definition is denoted by $\mathcal{Y}_{S\partial_{1}}$.
@@
    if $y^{\prime}$ is non-dominated and $y^{\prime}$ lies on the boundary of the upper image $\mathcal{Y}^{\geqq}\coloneqq\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}^{p}_{\geqq})$, i.e., $y^{\prime}\in\mathcal{Y}_{N}\cap\partial\mathcal{Y}^{\geq}$ (Eusébio and Figueira, 2009b, ; Eusébio and Figueira, 2009c, ; Liefooghe et al.,, 2014, 2015; Correia et al.,, 2021). The corresponding set of supported non-dominated points is denoted by $\mathcal{Y}_{S\partial_{2}}$.
@@
    if $y^{\prime}$ is nondomianted and $y^{\prime}$ is located on the non-dominated frontier defined as the set $\{y\in\operatorname{conv}(\mathcal{Y}_{N})\colon\operatorname{conv}(\mathcal{Y}_{N})\cap(y-\mathds{R}^{p}_{\geqq})=\{y\}\}$ (Hamacher et al.,, 2007), the set supported non-dominated points according to this definition is denoted by $\mathcal{Y}_{SNF}$.
  • (4)
    if $y$ is the image of a supported efficient solution, which are those efficient solutions that can be obtained as optimal solutions of a weighted sum scalarization with weights strictly greater zero, i.e. $y$ is an image of a solution $x\in\arg\min P_{\lambda}$ for a $\lambda\in\Lambda_{p}$ (Visẽe et al.,, 1998; Ehrgott,, 2005; Raith and Ehrgott,, 2009; Przybylski et al., 2010a, ; Argyris et al.,, 2011; Raith and Sedeño-Noda,, 2017). The set of supported non-dominated points obtained with respect to this definition is denoted by $\mathcal{Y}_{S\lambda}.$
Note, that some publications adopt the definition related to $\mathcal{Y}_{S\lambda}$ but allowing $\lambda\in\Lambda_{p}^{0}$, provided that $y$ is non-dominated, see, e.g. Gandibleux et al., (2001); Liefooghe et al., (2015). Regarding definition $\mathcal{Y}_{SNF}$amacher et al., (2007) uses the wording efficient frontier instead of the non-dominated frontier. There also exist definitions based on convex combinations of the non-dominated points as given in Özpeynirci and Köksalan, (2010).
It holds $\mathcal{Y}_{S\partial_{1}}=\mathcal{Y}_{S\partial_{2}}$ for general MOO problems after the definition, and from now on the definition $\mathcal{Y}_{S\partial}$ is used.
Theorem 3.2
$\mathcal{Y}_{S\partial_{1}}=\mathcal{Y}_{S\partial_{2}}$ for general MOO problems.
Proof.
Based on the known results $\mathcal{Y}_{N}=(\mathcal{Y}+\mathbb{R}^{p}_{\geqq})_{N}$ (Ehrgott,, 2005) and $(\text{conv}(\mathcal{Y}))_{N}=\text{conv}(\mathcal{Y}+\mathbb{R}^{p}_{\geqq})_{N}$(Benson,, 1998), it is easy to see that the following equalities hold:
$\mathcal{Y}_{S\partial_{1}}$ $=\partial(\text{conv}(\mathcal{Y}))\cap\mathcal{Y}_{N}=\partial(\text{conv}(\mathcal{Y}_{N}))\cap\mathcal{Y}_{N}$
$=\partial(\text{conv}(\mathcal{Y}_{N}+\mathbb{R}^{p}_{\geqq}))\cap\mathcal{Y}_{N}=\mathcal{Y}_{S\partial_{2}}.\qed$
@@
According to 2.4, any non-dominated point $y^{\prime}\in\mathcal{Y}$ in MOLP can be obtained as the image of an optimal solution $x\in\arg\min P_{\lambda}$ with $\lambda\in\Lambda_{p}$, and with 2.3 we obtain $\mathcal{Y}_{S\lambda}=\mathcal{Y}_{N}$. Any point $\hat{y}\in\mathcal{Y}_{S\partial}$ or $\hat{y}\in\mathcal{Y}_{SNF}$ is nondominated per definition and therefore $\hat{y}\in\mathcal{Y}_{S\lambda}.$ It follows
@@
Note that for an MOLP, $\mathcal{Y}$ is polyhedral and thus $\mathcal{Y}=\operatorname{conv}(\mathcal{Y})$. Consider $\hat{y}\in\mathcal{Y}_{S\lambda}$, i.e., $\hat{y}$ is an image of an optimal solution $x^{\prime}\in\arg\min P_{\lambda}$ with $\lambda\in\Lambda_{p}$. Since $\mathcal{Y}$ is polyhedral, $\hat{y}$ must lie in a face of $\operatorname{conv}(\mathcal{Y})$ (Nemhauser and Wolsey,, 1999). Hence, $\hat{y}\in\partial\operatorname{conv}(\mathcal{Y})$ and from 2.3 it follows $\hat{y}\in\mathcal{Y}_{N}$. Thus, it follows $\hat{y}\in\partial\operatorname{conv}(\mathcal{Y}_{N})\cap\mathcal{Y}_{N}$, and hence $\hat{y}\in\mathcal{Y}_{\partial}$. Furthermore, since $\operatorname{conv}(\mathcal{Y}_{N})$ is a nonempty convex set, it holds that $\hat{y}$ is a minimal element of $\operatorname{conv}(\mathcal{Y}_{N})$ induced by the cone $-\mathds{R}^{p}_{\geqq}$ (Boyd and Vandenberghe,, 2004) and we can conclude that $((\hat{y}-\mathds{R}^{p}_{\geqq})\backslash{\hat{y}})\cap\operatorname{conv}(\mathcal{Y}_{N})=\varnothing$. Consequently, it follows that $\hat{y}\in\mathcal{Y}_{SNF}$. Together it holds,
@@
However, this equivalence does not hold for the discrete or the non-linear case. In the literature, supported non-dominated points for MOCO are often characterized as non-dominated points on the boundary of the upper image and that they only lie on the maximally non-dominated faces, i.e., the non-dominated frontier. In contrast, the unsupported are characterized as non-dominated points that lie in the interior of the upper image, e.g., (Eusébio and Figueira, 2009c, ; Przybylski et al., 2010a, ; Correia et al.,, 2021).
Suppose a supported vector is defined according to Definition 3.1 (3) and (4), i.e., considering the sets $\mathcal{Y}_{S\lambda}$ and $\mathcal{Y}_{SNF}$. In that case, unsupported points may exist on the boundary of the upper image. Conversely, if a supported non-dominated vector is defined according to (1) and (2), i.e., points contained in $\mathcal{Y}_{S\partial}$, there may exist supported non-dominated points on the boundary of the upper image which are not lying on the non-dominated frontier, i. e., which lie on weakly non-dominated faces. This inconsistency motivates us to develop new, consistent definitions of supported and weakly supported non-dominated points.
@@
An efficient solution is called a weakly supported efficient solution if it is an optimal solution of a weighted-sum scalarization ${P}_{\lambda}$ for some weight $\lambda\in\Lambda_{d}^{0}$. Moreover, if the weight is strictly positive $\lambda\in\Lambda_{d}$ it is called supported efficient solution. The corresponding image is called weakly supported or supported (non-dominated) vector, respectively.
@@
  • $\mathcal{Y}_{wS}\subseteq\mathcal{Y}_{S\partial}$: Let $\hat{y}\in Y_{wS}$, i.e., there exists a $\lambda\in\Lambda^{0}_{p}$ such that $\hat{y}$ is an image of an optimal solution of ($P_{\lambda}$) and $\hat{y}$ is non-dominated.
    Suppose that $\hat{y}\notin\mathcal{Y}_{S\partial}$, i.e., $\hat{y}\notin\partial\operatorname{conv}(\mathcal{Y})$. Then $\hat{y}$ must lie in the interior of $\operatorname{conv}(\mathcal{Y})$. Thus, there exists an $\varepsilon$-neighborhood $B(\hat{y},\varepsilon)$ of $\hat{y}$, defined as $B(\hat{y},\varepsilon):=\hat{y}+B(0,\varepsilon)\subset\operatorname{conv}(\mathcal{Y})$, where $B(0,\varepsilon)$ is an open ball with radius $\varepsilon$ centered at the origin. Let $d\in\mathbb{R}^{p}_{>}$. Then we can choose some $\alpha\in\mathbb{R}$, $0<\alpha<\varepsilon$ such that $\alpha\,d\in B(0,\varepsilon)$. Now, $y^{\prime}=\hat{y}-\alpha\,d\in\operatorname{conv}(\mathcal{Y})$ with $y_{k}^{\prime}<\hat{y}_{k}$ for $k=1,\dots,p$ and
@@
  • $\mathcal{Y}_{S\partial}\subseteq\mathcal{Y}_{wS}:$ Let $\hat{y}\in\mathcal{Y}_{S\partial}\implies\hat{y}\in\partial\operatorname{conv}(\mathcal{Y})$ and $\hat{y}$ is non-dominated, hence it exists a supporting hyperplane $\{y\in\mathds{R}^{p}:\lambda^{\top}y=\lambda^{\top}\hat{y}\}$ with $\lambda\in\mathds{R}^{p}\setminus\{0\}$ and it holds $\lambda^{\top}y\geq\lambda^{\top}\hat{y}$ for all $y\in\operatorname{conv}(\mathcal{Y})$. Furthermore, by applying the seperation theorem to the dijoint sets $\{y+d\colon d\in\mathds{R}^{p}_{>}\}$ and $\{\hat{y}-d^{\prime}\colon d^{\prime}\in\mathds{R}^{p}_{>}\}$, we obtain:
@@
    for all $y\in\operatorname{conv}(\mathcal{Y})$, and $d,d^{\prime}\in\mathds{R}^{p}_{>}.$ Choose $d^{\prime}=e_{k}+\varepsilon\,e$, where $e_{k}$ is the $k$-th unit vector, $e=(1,\ldots,1)^{\top}\in\mathds{R}^{p}$. With $\varepsilon>0$ arbitrary small we see that $\lambda_{k}\geq 0$ for all $k=1,\ldots,p$. Thus, $\hat{y}$ is the image of an optimal solution to $P_{\lambda}$ with $\lambda\in\Lambda^{p}_{0}$ and is non-dominated. It follows $\hat{y}\in\mathcal{Y}_{wS}$.
@@
$\mathcal{Y}_{S}\subseteq\mathcal{Y}_{SNF}:$ Let $\hat{y}\in\mathcal{Y}_{S}=\mathcal{Y}_{S\lambda}$, i.e., $\hat{y}$ is the image of an optimal solution of $P_{\lambda}$ with $\lambda\in\Lambda_{p}$. Since $\operatorname{conv}(\mathcal{Y}_{N})$ is a nonempty convex set, $\hat{y}$ is a minimal element of $\operatorname{conv}(\mathcal{Y}_{N})$ induced of the cone $-\mathds{R}^{p}_{\geqq}$ and hence it holds $((\hat{y}-\mathds{R}^{p}_{\geqq})\backslash{\hat{y}})\cap\operatorname{conv}(\mathcal{Y}_{N})=\varnothing$ (Boyd and Vandenberghe,, 2004). Consequently, it follows that $\hat{y}\in\mathcal{Y}_{SNF}$.
$\mathcal{Y}_{S}\supseteq\mathcal{Y}_{SNF}:$ Any point $\hat{y}\in\mathcal{Y}_{SNF}$ is a minimal element of the set $\operatorname{conv}(Y_{N})$ induced by the cone $-\mathds{R}_{\geqq}^{p}$. Thus $\hat{y}$ is a nondominated point for the set $\operatorname{conv}(\mathcal{Y}_{N})$. If we define the relaxation MOLP of the given MOCO problem by
@@
it follows that $\hat{y}$, is an image of an optimal solution for the reduced MOLP. According to 2.4, there is a $\lambda\in\Lambda_{p}$ such that $\hat{y}$ is the image of an optimal solution to $P_{\lambda}$ and hence $\hat{y}\in\mathcal{Y}_{S}$. ∎
Consequently, Definitions 3.1 (1) and (2) include the weakly supported non-dominated points for general MOO problems, while Definitions (3) and (4) only contain the supported non-dominated points (in the discrete case). Hence, in the discrete case, the weakly supported non-dominated points lie on the boundary of the upper image while the supported ones lie only on the non-dominated frontier, i. e., on maximally non-dominated faces.
@@
To illustrate the geometrical properties of weakly supported non-dominated points, consider the outcome vectors in Figure 2. The outcome vectors are partitioned into two layers based on their value in the third component. The non-dominated points $y^{1}=(2,9,1),y^{2}=(3,6,1),y^{3}=(8,3,1),$ and $y^{4}=(6,5,1)$ all share a minimum value of $c_{3}=1$ in their third component (pink layer), while there are other non-dominated points having a value of 5 in the third component (blue layer).
@@
Let $\mathcal{Y}_{S}$ and $\mathcal{Y}_{wS}$ be the sets of supported and weakly supported non-dominated points of a MOCO problem, respectively. Then, $\mathcal{Y}_{S}\subseteq\mathcal{Y}_{wS}$ and there exist instances where $\mathcal{Y}_{S}\subset\mathcal{Y}_{wS}$, i. e., the set of supported non-dominated points is a proper subset of the set of the supported non-dominated points.
@@
The inclusion $\mathcal{Y}_{S}\subseteq\mathcal{Y}_{wS}$ holds per definition. Consider the outcome set $\mathcal{Y}=\{y^{1},y^{2},y^{3},y^{4}\}$ with $y^{1}=(2,9,1),y^{2}=(3,6,1),y^{3}=(8,3,1),$ and $y^{4}=(6,5,1)$, which represents the pink layer of the outcome set in Example 3.7, displayed in Figure 2. It is straightforward to construct an artificial MOCO problem with this outcome set.
All points $y^{i}$ with $i\in\{1,\ldots,4\}$ are weakly supported non-dominated points since their preimages are optimal solutions of a weighted sum problem $P_{\lambda}$ with $\lambda=(0,0,1)^{\top}$. Since $y^{1}$,$y^{2}$ and $y^{3}$ are also supported non-dominated points, since their preimages are optimal solutions of the respective weighted sum problems $P_{\lambda^{1}}$ with $\lambda^{1}=(0.7,0.1,0.2)^{\top}$, $P_{\lambda^{2}}$ with $\lambda^{2}=(0.4,0.4,0.2)^{\top}$, and $P_{\lambda^{3}}$ with $\lambda^{3}=(0.1,0.8,0.1)^{\top}$, respectively. The corresponding (projected) weight space decomposition is illustrated in Figure 3. Note that, for example, $(6,4.2)^{\top}\in\operatorname{conv}(\mathcal{Y}_{N})\cap(y^{4}-\mathds{R}^{p}_{\geqq})$ implying $y^{4}\notin\{y\in\operatorname{conv}(\mathcal{Y}_{N})\colon\operatorname{conv}(\mathcal{Y}_{N})\cap(y-\mathds{R}^{p}_{\geqq})=\{y\}\}$. By 3.6, it follows that $y_{4}\notin\mathcal{Y}_{S}$, although $y^{4}\in\mathcal{Y}_{wS}$.
@@
The clear distinction between the sets of supported non-dominated and weakly supported non-dominated solutions (Definition 3.4) is also necessary as the corresponding problems may differ in their output time complexity. For instance, in the case of the minimum cost flow problem, it can be shown that supported efficient solutions can be determined in output-polynomial time, while this is not the case for weakly supported solutions, unless $\mathbf{P}=\mathbf{NP}$ (Könen and Stiglmayr, 2023a, ).
@@
While the above characterization of supported points holds for MOCO problems, it does not extend to general MOO problems, and it can be shown that in the case of general MOO problems it may hold $\mathcal{Y}_{S}\subset\mathcal{Y}_{SNF}$. This distinction arises because the weighted sum scalarization for $\lambda\in\Lambda_{p}$ method determines only supported properly non-dominated points in the sense of Geoffrion (Geoffrion,, 1968), as discussed in Ehrgott, (2005). Note that for MOCO problems every efficient solution is also properly efficient (Ehrgott,, 2005). However, for general MOO problems, the set of properly supported non-dominated points may form a strict subset of the non-dominated points on the non-dominated frontier. The following example, similar to the ones in Boyd and Vandenberghe, (2004) or Ehrgott, (2005), illustrates this.
@@
Consider the points $y^{1}=(1,2)$ and $y^{2}=(2,1)$, which can be obtained with $\lambda^{1}=(0,1)$ and $\lambda^{2}=(1,0)$, respectively. It holds $y^{1},y^{2}\in\mathcal{Y}_{SNF}$ but $y^{1}$ and $y^{2}$ are not optimal solutions for $P_{\lambda}$ for any $\lambda\in\Lambda$ (Boyd and Vandenberghe,, 2004; Ehrgott,, 2005). Thus, $y^{1},y^{2}\notin\mathcal{Y}_{S}$. However, any point on the left lower boundary without the points $(1,2)$ and $(2,1)$ are both contained in $\mathcal{Y}_{SNF}$ and $\mathcal{Y}_{S}$. Therefore, in this example, $\mathcal{Y}_{S}\subset\mathcal{Y}_{SNF}$. The example is illustrated in Figure 4

### v2

@@ 2 Supported and Weakly Supported Non-dominated Points in MOCO problems
Several definitions and characterizations despite Definition˜2.1 of supported non-dominated points exist in the literature and sometimes are even inconsistent within a single publication.
@@
    if $y^{\prime}$ is non-dominated and $y^{\prime}$ lies on the boundary of the upper image $\mathcal{Y}^{\geqq}\coloneqq\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}^{p}_{\geqq})$, i.e., $y^{\prime}\in\mathcal{Y}_{N}\cap\partial\mathcal{Y}^{\geqq}$ (Eusébio and Figueira, 2009b, ; Eusébio and Figueira, 2009c, ; Liefooghe et al.,, 2014, 2015; Correia et al.,, 2021). The corresponding set of supported non-dominated points is denoted by $\mathcal{Y}_{S\partial}$.
@@
    if $y^{\prime}$ is non-dominated and $y^{\prime}$ is located on the non-dominated frontier defined as the set $\{y\in\operatorname{conv}(\mathcal{Y}_{N})\colon\operatorname{conv}(\mathcal{Y}_{N})\cap(y-\mathds{R}^{p}_{\geqq})=\{y\}\}$ (Hamacher et al.,, 2007), the set supported non-dominated points according to this definition is denoted by $\mathcal{Y}_{SNF}$.
@@
    if $y^{\prime}$ is the image of a supported efficient solution, which are those efficient solutions that can be obtained as optimal solutions of a weighted sum scalarization with weights strictly greater zero, i.e. $y$ is an image of a solution $x\in\arg\min P_{\lambda}$ for a $\lambda\in\Lambda_{p}$ (Visẽe et al.,, 1998; Ehrgott,, 2005; Raith and Ehrgott,, 2009; Przybylski et al., 2010a, ; Argyris et al.,, 2011; Raith and Sedeño-Noda,, 2017). The set of supported non-dominated points obtained with respect to this definition is denoted by $\mathcal{Y}_{S\lambda}.$
Note, that some publications adopt the definition related to $\mathcal{Y}_{S\lambda}$ but allowing $\lambda\in\Lambda_{p}^{0}$, provided that $y$ is non-dominated, see, e.g. Gandibleux et al., (2001); Liefooghe et al., (2015). Regarding definition $\mathcal{Y}_{SNF}$Hamacher et al., (2007) uses the wording efficient frontier instead of the non-dominated frontier. There also exist definitions based on convex combinations of the non-dominated points as given in Özpeynirci and Köksalan, (2010).
@@
According to ˜1.8, any non-dominated point $y^{\prime}\in\mathcal{Y}$ in MOLP can be obtained as the image of an optimal solution $x\in\arg\min P_{\lambda}$ with $\lambda\in\Lambda_{p}$, and with ˜1.5 we obtain $\mathcal{Y}_{S\lambda}=\mathcal{Y}_{N}$. Any point $\hat{y}\in\mathcal{Y}_{S\partial}$ or $\hat{y}\in\mathcal{Y}_{SNF}$ is non-dominated per definition and therefore $\hat{y}\in\mathcal{Y}_{S\lambda}.$ It follows
@@
[…]
@@
This equivalence does not hold for the discrete or the non-linear case. In the literature, supported non-dominated points for MOCO or MOILP are often characterized as non-dominated points that lie on the boundary of the upper image and that they only lie on the maximally non-dominated faces, i.e., the non-dominated frontier. In contrast, the unsupported are characterized as non-dominated points that lie in the interior of the upper image, e.g., (Eusébio and Figueira, 2009c, ; Przybylski et al., 2010a, ; Correia et al.,, 2021).
However, depending on the definition, unsupported non-dominated points may exist that lie on the boundary of the upper image or supported non-dominated points that lie on weakly non-dominated faces and thus do not lie on the non-dominated frontier. In this case, they cannot be obtained as optimal solutions of a weighted sum problem with weights strictly greater than zero. Suppose a supported non-dominated point is defined according to Definition˜2.1 (2) and (3), i.e., considering the sets $\mathcal{Y}_{S\lambda}$ and $\mathcal{Y}_{SNF}$. In that case, unsupported points may exist on the boundary of the upper image. Conversely, if a supported non-dominated point is defined according to (1), i.e., points contained in $\mathcal{Y}_{S\partial}$, there may exist supported non-dominated points on the boundary of the upper image which are not lying on the non-dominated frontier, i. e., which lie on weakly non-dominated faces. This inconsistency motivates us to develop new, consistent definitions of supported and weakly supported non-dominated points.
@@
An efficient solution is called a weakly supported efficient solution if it is an optimal solution of a weighted-sum scalarization ${P}_{\lambda}$ for some weight $\lambda\in\Lambda_{p}^{0}$. Moreover, if the weight is strictly positive $\lambda\in\Lambda_{p}$, it is called supported efficient solution. The corresponding image is called weakly supported or supported (non-dominated) point, respectively.
Note that the terminology weakly supported efficient solutions, despite the fact that these solutions are still efficient, is motivated by geometric considerations: any weakly supported non-dominated point that is not a supported non-dominated point lies on a weakly non-dominated face, but not on a non-dominated face. In contrast, supported points necessarily lie on non-dominated faces.
@@
With these definitions, the following characterizations for MOCO problems hold: While weakly supported non-dominated points lie on the boundary of the upper image, supported non-dominated points lie only on the non-dominated frontier, i. e., only on maximally non-dominated faces. The unsupported non-dominated points lie in the interior of the upper image.
@@
  • $\mathcal{Y}_{wS}\subseteq\mathcal{Y}_{S\partial}$: Let $\hat{y}\in\mathcal{Y}_{wS}$, i.e., there exists a $\lambda\in\Lambda^{0}_{p}$ such that $\hat{y}$ is an image of an optimal solution of ($P_{\lambda}$) and $\hat{y}$ is non-dominated.
    Suppose that $\hat{y}\notin\mathcal{Y}_{S\partial}$, i.e., $\hat{y}\notin\partial\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}_{\geqq}^{p})$. Then $\hat{y}$ must lie in the interior of $\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}_{\geqq}^{p})$. Thus, there exists an $\varepsilon$-neighborhood $B(\hat{y},\varepsilon)$ of $\hat{y}$, defined as $B(\hat{y},\varepsilon):=\hat{y}+B(0,\varepsilon)\subset\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}_{\geqq}^{p})$, where $B(0,\varepsilon)$ is an open ball with radius $\varepsilon$ centered at the origin. Let $d\in\mathbb{R}^{p}_{>}$. Then we can choose some $\alpha\in\mathbb{R}$, $0<\alpha<\varepsilon$ such that $\alpha\,d\in B(0,\varepsilon)$. Now, $y^{\prime}=\hat{y}-\alpha\,d\in\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}_{\geqq}^{p})$ with $y_{k}^{\prime}<\hat{y}_{k}$ for $k=1,\dots,p$ and
@@
  • $\mathcal{Y}_{S\partial}\subseteq\mathcal{Y}_{wS}:$ Let $\hat{y}\in\mathcal{Y}_{S\partial}\implies\hat{y}\in\partial\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}_{\geqq}^{p})$ and $\hat{y}$ is non-dominated, hence it exists a supporting hyperplane $\{y\in\mathds{R}^{p}:\lambda^{\top}y=\lambda^{\top}\hat{y}\}$ with $\lambda\in\mathds{R}^{p}\setminus\{0\}$ and it holds $\lambda^{\top}y\geq\lambda^{\top}\hat{y}$ for all $y\in\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}_{\geqq}^{p})$. Furthermore, by applying the separation theorem to the disjoint sets $\{y+d\colon d\in\mathds{R}^{p}_{>}\}$ and $\{\hat{y}-d^{\prime}\colon d^{\prime}\in\mathds{R}^{p}_{>}\}$, we obtain:
@@
    for all $y\in\operatorname{conv}(\mathcal{Y}_{N}+\mathds{R}_{\geqq}^{p})$, and $d,d^{\prime}\in\mathds{R}^{p}_{>}.$ Choose $d^{\prime}=e_{k}+\varepsilon\,e$, where $e_{k}$ is the $k$-th unit vector and $e=(1,\ldots,1)^{\top}\in\mathds{R}^{p}$. With $\varepsilon>0$ arbitrary small we see that $\lambda_{k}\geq 0$ for all $k=1,\ldots,p$. Thus, $\hat{y}$ is the image of an optimal solution to $P_{\lambda}$ with $\lambda\in\Lambda^{p}_{0}$ and per definition non-dominated. It follows $\hat{y}\in\mathcal{Y}_{wS}$.
Note that although the following results are stated for MOCO, they can be directly transferred to MOILP.
@@
[…]
$\mathcal{Y}_{S}\supseteq\mathcal{Y}_{SNF}:$ Any point $\hat{y}\in\mathcal{Y}_{SNF}$ is a minimal element of the set $\operatorname{conv}(\mathcal{Y}_{N})$ induced by the cone $\mathds{R}_{\geqq}^{p}$. Thus $\hat{y}$ is a non-dominated point for the set $\operatorname{conv}(\mathcal{Y}_{N})$. If we define the relaxation MOLP of the given MOCO problem by
@@
it follows that $\hat{y}$, is an image of an optimal solution for the reduced MOLP. According to ˜1.8, there is a $\lambda\in\Lambda_{p}$ such that $\hat{y}$ is the image of an optimal solution to $P_{\lambda}$ and hence $\hat{y}\in\mathcal{Y}_{S}$. ∎
Consequently, Definition 2.1 (1) includes the weakly supported non-dominated points for general MOO problems, while Definitions (3) and (4) only contain the supported non-dominated points (in the discrete case). Hence, in the discrete case, the weakly supported non-dominated points lie on the boundary of the upper image while the supported ones lie only on the non-dominated frontier, i. e., on maximally non-dominated faces.
@@
To illustrate the geometrical properties of weakly supported non-dominated points, consider the outcome vectors in Figure˜3. The outcome vectors are partitioned into two layers based on their value in the third component. The non-dominated points $y^{1}=(2,9,1),y^{2}=(3,6,1),y^{3}=(8,3,1),$ and $y^{4}=(6,5,1)$ all share a minimum value of $c_{3}=1$ in their third component (pink layer), while there are other non-dominated points having a value of 5 in the third component (blue layer).
@@
Let $\mathcal{Y}_{S}$ and $\mathcal{Y}_{wS}$ be the sets of supported and weakly supported non-dominated points of a MOCO problem, respectively. Then, $\mathcal{Y}_{S}\subseteq\mathcal{Y}_{wS}$ and there exist instances where $\mathcal{Y}_{S}\subset\mathcal{Y}_{wS}$, i. e., the set of supported non-dominated points is a proper subset of the set of the weakly supported non-dominated points.
@@
The inclusion $\mathcal{Y}_{S}\subseteq\mathcal{Y}_{wS}$ holds per definition. Consider the outcome set $\mathcal{Y}=\{y^{1},y^{2},y^{3},y^{4}\}$ with $y^{1}=(2,9,1),y^{2}=(3,6,1),y^{3}=(8,3,1),$ and $y^{4}=(6,5,1)$, which represents the pink layer of the outcome set in Example˜2.6, displayed in Figure 3. It is straightforward to construct an artificial MOILP problem with this outcome set.
All points $y^{i}$ with $i\in\{1,\ldots,4\}$ are weakly supported non-dominated points since their preimages are optimal solutions of a weighted sum problem $P_{\lambda}$ with $\lambda=(0,0,1)^{\top}$. Since $y^{1}$$y^{2}$ and $y^{3}$ are also supported non-dominated points, since their preimages are optimal solutions of the respective weighted sum problems $P_{\lambda^{1}}$ with $\lambda^{1}=(0.7,0.1,0.2)^{\top}$, $P_{\lambda^{2}}$ with $\lambda^{2}=(0.4,0.4,0.2)^{\top}$, and $P_{\lambda^{3}}$ with $\lambda^{3}=(0.1,0.8,0.1)^{\top}$, respectively. The corresponding (projected) weight space decomposition is illustrated in Figure˜4. Note that, for example, $(6,4.2,1)^{\top}\in\operatorname{conv}(\mathcal{Y}_{N})\cap(y^{4}-\mathds{R}^{p}_{\geqq})$ implies $y^{4}\notin\{y\in\operatorname{conv}(\mathcal{Y}_{N})\colon\operatorname{conv}(\mathcal{Y}_{N})\cap(y-\mathds{R}^{p}_{\geqq})=\{y\}\}$. By ˜2.5, it follows that $y_{4}\notin\mathcal{Y}_{S}$, although $y^{4}\in\mathcal{Y}_{wS}$.
@@
The clear distinction between the sets of supported non-dominated and weakly supported non-dominated solutions (Definition˜2.3) is also necessary as the corresponding problems may differ in their output time complexity. For instance, in the case of the minimum cost flow problem, it can be shown that supported efficient solutions can be determined in output-polynomial time, while this is not the case for weakly supported solutions, unless $\mathbf{P}=\mathbf{NP}$ (Könen and Stiglmayr,, 2025).
@@
While the above characterization of supported points holds for MOCO problems,it does not extend to general MOO problems, and it can be shown that in the case of general MOO problems it may hold $\mathcal{Y}_{S}\subset\mathcal{Y}_{SNF}$. This distinction arises because the weighted sum scalarization for $\lambda\in\Lambda_{p}$ method is only capable of identifying properly supported non-dominated points in the sense of Geoffrion (Geoffrion,, 1968), as discussed in Ehrgott, (2005). Note that for MOCO problems every efficient solution is also properly efficient (Ehrgott,, 2005). However, for general MOO problems, the set of properly supported non-dominated points may form a strict subset of the non-dominated points on the non-dominated frontier, highlighting the limitations of the weighted sum approach in capturing the entire set of supported solutions in general MOO problems (Ehrgott,, 2005). The following example, similar to the ones in Boyd and Vandenberghe, (2004) or Ehrgott, (2005), illustrates this.
@@
Consider the points $y^{1}=(1,2)$ and $y^{2}=(2,1)$, which can be obtained with $\lambda^{1}=(0,1)$ and $\lambda^{2}=(1,0)$, respectively. It holds $y^{1},y^{2}\in\mathcal{Y}_{SNF}$ but $y^{1}$ and $y^{2}$ are not optimal solutions for $P_{\lambda}$ for any $\lambda\in\Lambda$ (Boyd and Vandenberghe,, 2004; Ehrgott,, 2005). Thus, $y^{1},y^{2}\notin\mathcal{Y}_{S}$. However, any point on the left lower boundary without the points $(1,2)$ and $(2,1)$ are both contained in $\mathcal{Y}_{SNF}$ and $\mathcal{Y}_{S}$. Therefore, in this example, $\mathcal{Y}_{S}\subset\mathcal{Y}_{SNF}$. The example is illustrated in Figure˜5